(a) The Cartesian coordinates of a point in the xy-plane are (x, y) = (-3.44, -2.64) m. Find the polar coordinates of this point. r = m θ = ° (b) Convert (r, θ) = (4.73 m, 36.1°) to rectangular coordinates. x = m y = m
Answer #1

a)

radial co ordinate

r = sqrt ( 3.44^2 + 2.64^2) = 4.336

Angular co ordinate,

theta = 180 + arctan ( 2.64/ 3.44) = 217.5

(r, theta) = ( 4.336, 217.5)

========

b)

x = 4.73 cos 36.1 = 3.28

y = 4.73 sin 36.1 = 2.79

========

do comment in case any doubt, will reply for sure.. Goodluck